OpenStudy (anonymous):

Which rule will find the nth term of an arithmetic sequence with a rate of increase equal to 1/2 and a6 = 18? A. an = 30 + B. an = C. an = 15 − D. an = 15 +

OpenStudy (mathmate):

I believe all the responses should read: A. an=30+n/2 B. an=n/2 C. an=15-n/2 D. an=15+n/2

OpenStudy (anonymous):

yes thank you , you would be correct haha i messed up. but do you have the awnser please?

OpenStudy (amistre64):

why dont you just use n=6 to see if you get 18 as a result now?

OpenStudy (anonymous):

im sorry , i dont understand..

OpenStudy (amistre64):

thats rather sad :/ replace the "n" with the value of "6" and work the math .... if you end up with a6 = 18 then youve got the correct option.

OpenStudy (anonymous):

I think i got it ! thanks ! i first did it and thought it was A then i think i realised i did the problem in the wrong order cause you solve the division first so then i got D . I think the awnser is D

OpenStudy (mathmate):

Yes, you have the correct answer. Hope you worked out the way @amistre64 suggested.

OpenStudy (amistre64):

15 + 6/2 = 18, i like D as well :)

OpenStudy (anonymous):

Thank you! (: